# Decoupling the “What” and “Where” With Polar Coordinate Positional Embeddings

### v1

@@ Abstract
The attention mechanism in a Transformer architecture matches key to query based on both content—the what—and position in a sequence—the where. We present an analysis indicating that what and where are entangled in the popular RoPE rotary position embedding. This entanglement can impair performance particularly when decisions require independent matches on these two factors. We propose an improvement to RoPE, which we call Polar Coordinate Position Embeddings or PoPE, that eliminates the what-where confound. PoPE is far superior on a diagnostic task requiring indexing solely by position or by content. On autoregressive sequence modeling in music, genomic, and natural language domains, Transformers using PoPE as the positional encoding scheme outperform baselines using RoPE with respect to evaluation loss (perplexity) and downstream task performance. On language modeling, these gains persist across model scale, from 124M to 774M parameters. Crucially, PoPE shows strong zero-shot length extrapolation capabilities, whereas RoPE’s performance degrades significantly on longer sequences at test time without fine tuning or the use of position-interpolation methods.
@@ 1 Introduction
Two approaches were common: recurrent neural nets (RNNs) and slot-based encodings. RNNs are a means of encoding sequence position implicitly in the ordering of inputs. By contrast, slot-based encodings partition an input vector into separate components for each sequence step. A significant advantage of RNNs is their approximate translation equivariance, meaning that shifting the absolute position of a sub-sequence in the input yields analogous shifts in the model state. Slot-based representations do not share this property: learning to respond to content in one slot does not generalize to the same content in other slots.
@@
In this article, we argue that the Rotary Position Embedding (RoPE) (Su et al., 2024), a widely adopted solution, entangles the what and where in a way that can impair model performance, particularly when decision making requires independent matches on these two factors. We propose an alternative technique, which we call PoPE, that shares the advantages of RoPE over other positional encodings while allowing the Transformer to implement rules in which the key-query match can be characterized as a conjunction of a what match and a where match. Although PoPE is only a minor modification of RoPE, it introduces a powerful inductive bias that improves data efficiency, asymptotic accuracy, and yields superior context-length generalization.
@@ 4 Results
We introduce a task that requires identifying a target character within a variable-length source string. The target is defined to be at a specified relative offset (left or right) from a specified source character. For example, in the input QEOHoUbKfeSrMVNlCzXu, z, -3, N; the target N is three places to the left of source z in the string. Predicting the final (target) character of this sequence requires models to learn to independently manipulate the content and positional information of tokens and to apply pointer arithmetic operations.
@@
The dataset is constructed by procedurally generating examples of source strings, source symbols, and relative shifts. We compare RoPE (Su et al., 2024) against PoPE by training two Transformer models with cross-entropy loss applied only on the final (target) token and evaluated on the accuracy of final token. Table 1 shows the mean and standard deviation (3 seeds) in final token accuracy on the test set. While RoPE struggles to learn this task and reaches an average accuracy of just 11%, our method PoPE solves the task almost perfectly, reaching an average accuracy of nearly 95%. This result highlights the difficulty that RoPE has in teasing apart ‘what’ and ‘where’ information, i.e., identifying the position of certain content and determining the content at a certain relative position. In contrast, PoPE efficiently learns a generalizable solution for the task, presumably because ’what’ and ’where’ information can be disentangled in key-query matching.
@@
Figure 2: Length extrapolation at test-time on PG-19 dataset for different model sizes. We evaluate the models using RoPE (red) and PoPE (green) without any fine-tuning or position interpolation techniques at test-time on longer sequences (multiples of 1024 up to 10240).
@@
We measure the ability of PoPE to generalize to test-time sequences that are longer than those presented during training. We examine models pretrained on OpenWebText using a sequence length (context window) of 1024 tokens and assess zero-shot perplexity on much longer sequences (up to 10240 tokens) from the test split of the PG-19 dataset (Rae et al., 2020). PoPE shows strong out-of-the-box length extrapolation capabilities without any fine-tuning or position interpolation techniques at test time on 10x longer sequences (Figure 2). In stark contrast, RoPE’s performance significantly degrades on longer sequences at test time. It is also interesting to note that RoPE’s extrapolation performance degrades with model size, PoPE’s extrapolation performance remains largely stable. RoPE’s failure is due to its allowance of a what-where interaction: aspects of the key and query representation can dynamically shift the position tuning of a component. These shifts become problematic, especially for the lowest frequency components, as they exert their influence only when the context window is expanded.
@@ 5 Related Work
The original Transformer (Vaswani et al., 2017) uses sinusoidal embeddings to encode absolute positions; which are added to the token embeddings only at the input layer. The sinusoidal positional embeddings also generally underperform w.r.t more modern methods for relative positional embeddings (Su et al., 2024) which inject this information at every layer. It has been shown that autoregressive Transformers do not require explicit encoding of positional information to operate (Irie et al., 2019; Irie, 2025). This is also somewhat reminiscent of fast-weight programmers (FWPs) (Schmidhuber, 1992), which are now known as unnormalized linear Transformers that also did not use explicit positional encoding. Such Transformers tend to have better length extrapolation properties than both absolute and relative positional embeddings (Kazemnejad et al., 2023), although this comes at the expense of their in-distribution performance. Shaw et al. (2018) introduced a relative positional embedding which uses a separate set of keys that are selected based on the distance of the key and query. Music Transformer (Huang et al., 2019) proposed a more efficient implementation of this relative positional embedding. Dai et al. (2019) propose a different variant, where instead of learning a separate key for each offset, they use a learned projection of sinusoidal “offset encodings”, which are inspired by the absolute positional encodings. This approach generalizes better to longer lengths in practice. Additionally, the authors train sequentially within documents and allow attention to the previous batch, enhancing the long-range capabilities of the models further. Wang et al. (2020) generalized this idea of “offset encodings” by using complex-valued embeddings of inputs to encode information about the content, global position and their order relationships within the sequence. T5 (Raffel et al., 2020) takes a different approach: their method groups token pairs in log-sized buckets based on their distance, and it adds a learned per-bucket bias to the attention logit, allowing it to decrease the importance of far-away context. ALiBi Press et al. (2022) takes inspiration from this, adding learned scores to the attention matrix that decay with relative distance. Geometric (Csordás et al., 2022) or stick-breaking attention (Tan et al., 2025) take a radically different approach: it replaces the softmax activation function with a stick-breaking process that gives priority to good matches that are nearby without explicitly encoding positions or offsets. The authors claim superior length generalization.

### v2

@@ Abstract
The attention mechanism in a Transformer architecture matches key to query based on both content—the what—and position in a sequence—the where. We present an analysis indicating that what and where are entangled in the popular rotary position embedding (RoPE). This entanglement can impair performance particularly when decisions require independent matches on these two factors. We propose an improvement to RoPE, which we call Polar Coordinate Position Embedding or PoPE, that eliminates the what-where confound. PoPE is far superior on a diagnostic task requiring indexing solely by position or by content. On autoregressive sequence modeling in music, genomic, and natural language domains, Transformers using PoPE as the positional encoding scheme outperform baselines using RoPE with respect to evaluation loss (perplexity) and downstream task performance. On language modeling, these gains persist across model scale, from 124M to 774M parameters. Crucially, PoPE shows strong zero-shot length extrapolation capabilities compared not only to RoPE but even a method designed for extrapolation, YaRN, which requires additional fine tuning and frequency interpolation.
@@ 1 Introduction
In the 1980s, two approaches were common: recurrent neural nets (RNNs) and slot-based encodings. RNNs are a means of encoding sequence position implicitly in the ordering of inputs. By contrast, slot-based encodings partition an input vector into separate components for each sequence step. A significant advantage of RNNs is their approximate translation equivariance, meaning that shifting the absolute position of a sub-sequence in the input yields analogous shifts in the model state. Slot-based representations do not share this property: learning to respond to content in one slot does not generalize to the same content in other slots.
@@
In this article, we argue that the Rotary Position Embedding (RoPE) (Su et al., 2024), a widely adopted solution, entangles the what and where in a way that can impair model performance, particularly when decision making requires independent matches on these two factors. We propose an alternative technique, which we call PoPE, that shares the advantages of RoPE over other positional encodings while allowing the Transformer to implement rules in which the key-query match can be characterized as a conjunction of a what match and a where match. Although PoPE is only a minor modification of RoPE, it introduces a powerful inductive bias that improves data efficiency, asymptotic accuracy, and yields context-length generalization superior to state-of-the-art methods.
@@ 4 Results
We introduce a task that requires identifying a target character within a variable-length source string. The target is defined to be at a specified relative offset (left or right) from a specified source character. For example, in the input QEOHoUbKfeSrMVNlCzXu, z, -3, N; the target N is three places to the left of source z in the string. Predicting the final (target) character of this sequence requires models to learn to independently manipulate the content and positional information of tokens and to apply pointer arithmetic operations. The dataset is constructed by procedurally generating examples of source strings, source symbols, and relative shifts. We compare RoPE (Su et al., 2024) against PoPE by training two Transformer models with cross-entropy loss applied only on the final (target) token and evaluated on the accuracy of final token. Table 1 shows the mean and standard deviation (3 seeds) in final token accuracy on the test set.
@@
While RoPE struggles to learn this task and reaches an average accuracy of just 11%, our method PoPE solves the task almost perfectly, reaching an average accuracy of nearly 95%. This result highlights the difficulty that RoPE has in teasing apart ‘what’ and ‘where’ information, i.e., identifying the position of certain content and determining the content at a certain relative position. In contrast, PoPE efficiently learns a generalizable solution for the task, presumably because ’what’ and ’where’ information can be disentangled in key-query matching.
@@
Figure 2: Length extrapolation at test-time on PG-19 dataset for different model sizes. We evaluate baselines that use RoPE (red) or YaRN (yellow) against PoPE (green) which does not apply any fine-tuning or interpolation techniques and PoPE+ft (blue) which only uses fine-tuning. Sequences at test-time are multiples of 1024 up to 10240.
@@
We measure the ability of PoPE to generalize to test-time sequences that are longer than those presented during training. We examine models pretrained on OpenWebText using a sequence length (context window) of 1024 tokens and assess zero-shot perplexity on much longer sequences (up to 10240 tokens) from the test split of the PG-19 dataset (Rae et al., 2020). We also compare against a strong baseline, YaRN (Peng et al., 2024), which is a state-of-the-art method used to improve the length extrapolation capabilities of RoPE. YaRN applies an interpolation scheme to the base frequencies $\theta_{c}$ of RoPE and finetunes the model on longer sequences. For PoPE, we simply finetune on longer sequences without interpolating the frequency components, we refer to this variant as PoPE+ft in Figure 2. Both YaRN and PoPE+ft models are finetuned on sequences of length 4096 from the OpenWebText dataset for 500 steps with a warmup of 20 steps using batch size of 64 and learning rate of 6e-5 without any decay. In Figure 2, we see that RoPE’s performance significantly degrades on longer sequences at test time. YaRN mitigates this performance degradation on sequences whose length does not exceed the size of the context window used for fine-tuning (upto 4096). However, when extending to test sequences whose length exceeds that used for fine-tuning, YaRN’s performance also degrades significantly. In stark contrast, PoPE shows strong out-of-the-box length extrapolation capabilities without any fine-tuning or position interpolation techniques at test time on 10x longer sequences (Figure 2) and even outperforms specialized baselines like YaRN. Further, our fine-tuned PoPE variant, shows noticeable improvement in perplexity on longer sequences compared to PoPE, which suggests that the fine-tuning process enables the adaptation of low frequency components based on the longer sequences seen during fine-tuning. Overall, these results highlight the strengths of our simple method to overcome these critical issues of RoPE in a principled way. It is also interesting to note that RoPE’s extrapolation performance degrades with model size, while PoPE’s extrapolation performance remains largely stable. RoPE’s failure is due to its allowance of a what-where interaction: aspects of the key and query representation can dynamically shift the position tuning of a component. These shifts become problematic, especially for the lowest frequency components, as they exert their influence only when the context window is expanded.
@@ 5 Related Work
The original Transformer (Vaswani et al., 2017) uses sinusoidal embeddings to encode absolute positions; which are added to the token embeddings only at the input layer. The sinusoidal positional embeddings also generally underperform w.r.t more modern methods for relative positional embeddings (Su et al., 2024) which inject this information at every layer. It has been shown that autoregressive Transformers do not require explicit encoding of positional information to operate (Schmidhuber, 1992a; Irie et al., 2019; Irie, 2025). Such Transformers tend to have better length extrapolation properties than both absolute and relative positional embeddings (Kazemnejad et al., 2023), although this comes at the expense of their in-distribution performance. In the 1990s, neural sequence models, e.g., the Neural History Compressor (Schmidhuber, 1992b), used a relative positional encoding based on the inverse time that went by since the last unexpected input. Shaw et al. (2018) introduced a relative positional embedding which uses a separate set of keys that are selected based on the distance of the key and query. Music Transformer (Huang et al., 2019) proposed a more efficient implementation of this relative positional embedding. Dai et al. (2019) propose a different variant, where instead of learning a separate key for each offset, they use a learned projection of sinusoidal “offset encodings”, which are inspired by the absolute positional encodings. This approach generalizes better to longer lengths in practice. Additionally, the authors train sequentially within documents and allow attention to the previous batch, enhancing the long-range capabilities of the models further. Wang et al. (2020) generalized this idea of “offset encodings” by using complex-valued embeddings of inputs to encode information about the content, global position and their order relationships within the sequence. T5 (Raffel et al., 2020) takes a different approach: their method groups token pairs in log-sized buckets based on their distance, and it adds a learned per-bucket bias to the attention logit, allowing it to decrease the importance of far-away context. ALiBi Press et al. (2022) takes inspiration from this, adding learned scores to the attention matrix that decay with relative distance. Geometric (Csordás et al., 2022) or stick-breaking attention (Tan et al., 2025) take a radically different approach: it replaces the softmax activation function with a stick-breaking process that gives priority to good matches that are nearby without explicitly encoding positions or offsets. The authors claim superior length generalization.
@@ 6 Conclusion
7 Ethics Statement
We consider our work to be fundamental research on sequence modeling with Transformers with no direct societal implications. However, our work which develops a novel relative positional encoding scheme that improves sequence modeling capabilities in Transformer models could potentially lead to benefits as well as unforeseen harms and risks from dual-use applications. By improving sequence modeling capabilities across domains such as music, genomics and natural language, this work could potentially enhance the performance of Foundation Models used for downstream applications such as protein folding prediction, drug discovery, music composition and conversational assistance. The enhanced length extrapolation capabilities of our method are particularly valuable for real-world deployment of large language models, especially variants that use chain-of-thought as they process much longer sequences at test-time, potentially reducing the amount of fine tuning on longer sequences and the associated computational costs. However, as with any advancement in the capabilities of Foundation Models, it could potentially be misused for generating harmful content at scale or creating more sophisticated deepfakes in text, audio, images etc.
@@ Appendix C Additional Visualizations
Visualization of the learnable bias.
Below we visualize the learned biases $\delta_{c}$ from the pretrained PoPE models on OpenWebText from Table 4. Note that the biases were initialized as zeros and are restricted to lie in the range $2\pi$ to 0. The one consistent pattern across both model sizes is that the bias values ($\delta_{c}$) tend to deviate away from zero only for the higher frequency components, while the low frequency components remain at their initial zero values.
Figure 7: Visualization of the learned biases for the 124M pretrained PoPE model across all layers.
Figure 8: Visualization of the learned biases for the 253M pretrained PoPE model across all layers.